Shift report 20011124 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 24 2001 - 15:51:21 EST

  • Next message: Apache: "Shift report 20011125 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Nov 24 15:51:21 2001, Shift: 08:00-16:00
    Supervisor: rw, with: 
    Log book#11 pages 31 - 31
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam for the entire shift. Booster problem from the previous 
    shift was fixed at 9:30. At 10:30 AGS BTA magnet power supply
    problem was announced. No time estimate for beam.
    
    2.run 5075, 5976,5977 are pedestal and DVM runs.
    
    3. No detector problems.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 24 2001 - 15:52:27 EST